OpenStudy (anonymous):

solve the equation. check the solution. -2/x+4 = 4/x+3

OpenStudy (anonymous):

\[\frac{-2}{x+4}=\frac{4}{x+3}\] \[-2(x+3)=4(x+4)\] etc

OpenStudy (anonymous):

so x= -11/3 ?